the car drove 3 km to the north and 4 km to the east. The distance covered by the car and it's displacement are:



A) 5km and 7km

B) 7 km and 1 km

c) 5 km and 1 km

7) 7 km and 5 km


1
Expert's answer
2021-05-03T10:33:50-0400

"d=3+4=7\ km\\s=\sqrt{3^2+4^2}=5\ km"

Answer: d) 7 km and 5 km